# Diffusion MRI with Machine Learning

### v3

@@ V-B2 Unpredictable generalizability to unseen test data
Generalizability of machine learning methods to new test data can be affected by factors that may not be clear beforehand [168]. For instance, it has been suggested that factors such as signal intensity and acquisition parameters such as echo time may impact the performance of a neural network model to the extent that a complete re-training may become necessary [371, 16]. Performance of conventional (i.e., non-learning) methods is probably much less affected by such unforeseen factors.
Closely related to the above point is the issue of out-of-distribution (OOD) data, which refers to data samples that are outside the distribution of the training data. Detecting OOD data in higher-dimensional signal spaces is very challenging, and machine learning methods have no performance guarantee on OOD data [372, 373]. A machine learning technique that works well on in-distribution data can experience unexpected and catastrophic failure on OOD test data. Obtaining training data that is sufficiently rich and heterogeneous to represent all data that will be potentially encountered at test time can be difficult if not impossible because many factors may influence the data distribution. In tractography, for example, data noise and artifacts can be important factors and the trained model may produce inferior results if the training and test data are different with regard to these aspects [18, 238]. Non-learning methods are less prone to such failures.
It has been shown that machine learning methods can lead to systematic and unpredictable biases in microstructure mapping from dMRI signal [147, 168]. Moreover, this bias can depend on factors such as SNR, partial volume effects, and distribution of the microstructure values in the training and test data [147, 168]. As a result, predictions of these methods may lead to erroneous conclusions when studying the impact of pathology on microstructure [168]. While such biases may be present in many machine learning methods, discovering them requires extensive validation and analysis that is missing in most published works. These suggest that machine learning methods may suffer from shortcomings that are not easy to recognize. In general, most modern machine learning methods are complex and difficult to interpret, a topic that is further discussed below in Section V-E.

### v5

@@ V-B2 Unpredictable generalizability to unseen test data
Generalizability of machine learning methods to new test data can be affected by factors that may not be clear beforehand [168]. For instance, it has been suggested that factors such as signal intensity and acquisition parameters such as echo time may impact the performance of a neural network model to the extent that a complete re-training may become necessary [372, 16]. Performance of conventional (i.e., non-learning) methods is likely much less affected by such unforeseen factors.
Closely related to the above point is the issue of out-of-distribution (OOD) data, which refers to data samples that are outside the distribution of the training data. Detecting OOD data in higher-dimensional signal spaces is very challenging, and machine learning methods have no performance guarantee on OOD data [373, 374]. A machine learning technique that works well on in-distribution data can experience unexpected and catastrophic failure on OOD test data. Obtaining a training dataset that is sufficiently rich and heterogeneous to represent all data that will be potentially encountered at test time can be difficult if not impossible because many factors may influence the data distribution. In tractography, for example, data noise and artifacts can be important factors and the trained model may produce inferior results if the training and test data are different with regard to these aspects [18, 238]. Non-learning methods are less prone to such failures.
It has been shown that machine learning methods can lead to systematic and unpredictable biases in microstructure mapping from dMRI signal [147, 168]. Moreover, this bias can depend on factors such as SNR, partial volume effects, and distribution of the microstructure values in the training and test data [147, 168]. As a result, predictions of these methods may lead to erroneous conclusions when studying the impact of pathology on microstructure [168]. While such biases may be present in many machine learning methods, one will require extensive validation and analysis to discover those biases. Most published works lack adequate experiments to address this issue. These suggest that machine learning methods may suffer from shortcomings that are not easy to recognize. In general, most modern machine learning methods are complex and difficult to interpret, a topic that is further discussed below in Section V-E.
@@ VI Conclusions
Declaration of Competing Interests
The author has no conflicts of interest to declare.
Data and Code Availability
This paper does not include any data or code.
Author Contributions
Davood Karimi: Conceptualization; search and review of the literature; formal analysis; Original draft preparation; review and editing; funding acquisition.